Second problem: (5x) times (5x) (to the third)

Answer by sdmmadam@yahoo.com(530)

Directions on little brother's lesson is: Write each expression using a single exponet. Problem is:
1)(5x) (to the 2nd) times x (to the 6th) times x (to the third)
=[(5x)^2]X (x^6) X (x^3)
=25(x^2)(x^6)(x^3)
=25[(x)^(2+6+3)] (when quantities of the same base are multiplied write the base and add the powers)
=25 (x)^11
Second problem: (5x) times (5x) (to the third)
2)(5x) times (5x) (to the third)
=(5x)X[(5x)^3]
=[(5x)^(1+3)] (when quantities of the same base are multiplied write the base and add the powers)
=(5x)^4
=(5^4)(x^4)
=625(x^4)